OpenStudy (anonymous):

how do you do 1/10 + 3/5 ?

OpenStudy (anonymous):

0.7

OpenStudy (sriram):

LCM=10 3/5=6/10 hence add now

OpenStudy (anonymous):

but if i add it would be 9/15 wouldnt it?

OpenStudy (sriram):

NO u dont add the denomiNATORS LIKE 1/2+1/2=HALF +HALF=1 and not=2/4

OpenStudy (anonymous):

i dont get it ;/

OpenStudy (mathhelp346):